OpenStudy (anonymous):

find a formula for the sum of n terms. Use the formula to find the limit as n->infinity limit as n goes to infinity of the sum when i = 1 (1+i/n)(2/n)

OpenStudy (anonymous):

The equation looks more so like this \[\lim_{n \rightarrow \infty} \sum_{i=1}^{n} (1+\frac{ i }{ n })(\frac{ 2 }{ n })\]

OpenStudy (anonymous):

I got 3

OpenStudy (anonymous):

I know that is the answer but I am having trouble starting it. What could I take out to get just 1+ i on the inside?

OpenStudy (anonymous):

you can take out (2/n) since it's *i* that you're incrementing on

OpenStudy (anonymous):

so I would have (1+i)(4/n^2) ?

OpenStudy (anonymous):

\[(2/n)\sum_{i=1}^{n}(1+i/n) = (2/n) (\sum_{i=1}^{n}1 + \sum_{i=1}^{n}i)\]

OpenStudy (anonymous):

the first sum is just n the second sum is (n/2)(n+1)

OpenStudy (anonymous):

ahhhh yes thanks!

OpenStudy (anonymous):

now you just need to multiply them all together and take the limit. Answer should be 3

OpenStudy (anonymous):

Yea i checked back of the book for final forumula and limit so, anyways thanks all I needed was the starting step! :)

OpenStudy (anonymous):

opps the second sum should be \[(1/n)\sum_{i=1}^{n}i\]

OpenStudy (anonymous):

right now i got \[\lim_{n \rightarrow \infty} \frac{ 2 }{ n } (n) + \frac{ 1 }{ n } (\frac{ n^2+n }{ 2 }\] and i multiplied the outsides of the 2 to get \[\lim_{n \rightarrow \infty} \frac{ 2n }{ n } + \frac{ n^2+n }{ 2 }\] is that correct?

OpenStudy (anonymous):

nope, you probably messed your algebra somewhere

OpenStudy (anonymous):

there is no n^2 term

OpenStudy (anonymous):

the sum of i though is \[\frac{ n(n+1) }{ 2 }\] is it not?

OpenStudy (anonymous):

yes but there is also (1/n) in front of the sum. I believe I corrected mistake above

OpenStudy (anonymous):

ahh so that neutralizes the n in front

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so if I simplify I end up with \[\lim_{n \rightarrow \infty} 2 + \frac{ n+1 }{ }\] right?

OpenStudy (anonymous):

woops the n+1 is over 2

OpenStudy (anonymous):

still no. Did add the second sum to the first sum?

OpenStudy (anonymous):

you should have (2/n) [n + (n+1)/2]

OpenStudy (anonymous):

ok then the end is \[\lim_{n \rightarrow \infty} \frac{ 1 }{ n } (\frac{ 3n+1 }{ 1 })\] and the limit is 3

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

thanks!